OpenStudy (thomasfire55):

Rewrite the following quadratic function in vertex form. Then, determine the axis of symmetry. y = 5x^2 + 15x - 2

OpenStudy (cosmo):

\[y=5x ^{2}+15x-2\] \[y=5(x^2+3)-2\] \[(3/2)^2=9/4\] \[y=5(x+1.5)^2-2-(9/4)\]

jhonyy9 (jhonyy9):

inside parantheses next 3 you missed an x cosmo !

jhonyy9 (jhonyy9):

@Cosmo ATTENTION !!!

OpenStudy (cosmo):

y=5x2+15x−2 y=5(x2+3x)−2 (3/2)2=9/4 y=5(x+1.5)2−2−(9/4)